# Control of kinetic opinion dynamics in popularity-adaptive social networks

### v1

@@ 4.3 Test 3: emergence of echo chambers under local contact control
In this test we analyse a setting in which two groups of leaders, characterised by markedly opposite opinions, interact with a large population of followers whose initial opinions are closer to one of the two leader groups. The aim is to study how local control of contacts performed by one or both leader groups affects the formation of echo chambers and the eventual opinion distribution of the mass.
The total population consists of interacting agents, with $25\%$ assigned to Leader Group A, $25\%$ to Leader Group B, and the remaining $50\%$ form the follower population. The simulation horizon is $T=150$ with time step $\Delta t=10^{-3}$.
@@
The dynamics of contacts and opinions evolve according to the parameters reported in Table 1, here with penalization weights $\gamma_{c}=\gamma_{v}=1$.
The initial joint density $f_{0}(c,v)$, sampled from $N_{s}=10^{6}$ agents, is illustrated in Fig. 9. At time $t=0$, leaders of Group A are sampled from $(c,v)\sim\mathrm{Unif}\big([200,250]\times[-0.6,-0.4]\big),$ and are associated with the target opinion $\tilde{v}_{A}=-0.5$, representing a highly connected community endorsing a moderately negative stance. Leaders of Group B are initialised as $(c,v)\sim\mathrm{Unif}\big([200,250]\times[0.4,0.6]\big),$ with a corresponding positive target opinion $\tilde{v}_{B}=0.5$. Followers are drawn from $(c,v)\sim\mathrm{Unif}\big([50,100]\times[0.1,0.6]\big),$ representing a poorly connected and more heterogeneous population.
@@
The results, displayed in Fig. 10 through snapshots of the joint density $f(c,v,t)$, and in Fig. 12 via the evolution of the marginal distributions of contacts $h(c)$ and opinions $g(v)$, reveal a clear differentiation among the three scenarios. In all cases, both leader groups initially experience a decline in their number of contacts due to the endogenous mechanisms governing the contact dynamics. When only the opinion control is active, this decay leads to a strong weakening of both leader groups, and the followers tend to preserve their initial opinion distribution. When Group A activates the contact control, however, the decline of its connectivity is mitigated and a cohesive cluster emerges around its target opinion; this cluster progressively attracts both the follower population and, indirectly, the weakened Group B, whose loss of contacts reduces its influence. In the third scenario, where both leader groups activate the contact control, two persistent and well-connected leader clusters form around their respective target opinions, preventing their decay. In this case, a portion of the follower population converges towards an intermediate compromise opinion, remaining influenced by the simultaneous presence of the two stable leader communities: in other words, a fraction is pulled toward each pole, producing a stable bi-polarization or two echo chambers with a residual central compromise group.

### v2

@@ 4.3 Test 3: emergence of echo chambers under local contact control
In this test, we analyze a setting in which two groups of leaders, characterized by markedly opposite opinions, interact with a large population of followers whose initial opinions are closer to one of the two leader groups. The aim is to study how local control of contacts performed by one or both leader groups affects the formation of echo chambers and the eventual opinion distribution of the mass.
We consider three population of interacting agents, with $25\%$ assigned to Leader Group A ($\ell=A$), $25\%$ to Leader Group B ($\ell=B$), and the remaining $50\%$ form the follower population ($\ell=F$). The simulation horizon is $T=150$ with time step $\Delta t=10^{-3}$.
@@
The dynamics of contacts and opinions evolve according to the parameters reported in Table 1, here with penalization weights $\gamma_{c}=\gamma_{v}=1$ for both populations of leaders.
The total initial joint density $f_{0}(c,v)$, sampled from $N_{s}=10^{6}$ agents, is illustrated in Fig. 10. At time $t=0$, leaders of Group A are sampled from $(c,v)\sim\mathrm{Unif}\big([200,250]\times[-0.6,-0.4]\big),$ and are associated with the target opinion $\tilde{v}_{A}=-0.5$, representing a highly connected community endorsing a moderately negative stance. Leaders of Group B are initialised as $(c,v)\sim\mathrm{Unif}\big([200,250]\times[0.4,0.6]\big),$ with a corresponding positive target opinion $\tilde{v}_{B}=0.5$. Followers are drawn from $(c,v)\sim\mathrm{Unif}\big([50,100]\times[0.1,0.6]\big),$ representing a poorly connected and more heterogeneous population.
@@
[…]